

********************************************************************
UARS WEEKLY STATUS REPORT - 02 AUG 2002
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
Nothing received.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
PEM continues to work with the GUVI team on the diagnosis of data from
the first conjunction period. Analysis of the two conjunction passes showed
that the aurora as detected by GUVI was at higher latitude than UARS.
GUVI images from other times from this conjunction period show spatial
regions that both GUVI and PEM sampled when aurora was present. Separation
is several minutes to 10's of minutes.
PEM data from the second GUVI-PEM conjunction is now being processed.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered. NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations to support re-engineering of the UARS
Mission Operations Concept to develop capabilities for the
Science Traceability Mission.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
***********
FOT
***********
REPORTING PERIOD
This report covers 20 July 2002 (Orbit 59362, GMT Day 201) through 26
July 2002 (Orbit 59467, GMT Day 207).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
WINDII 202/0056 59375 POWER OFF
HALOE 202/1335 59385 POWER ON
The beta angle ranged between the angles of 53.9 degrees to 28.1
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 02 August 2002 (DOY 214).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.29 to 1.34 degrees C. Additional
battery performance data for each battery on 19 July and 26 July 2002
are provided in Appendix D.
The clock error ranged from 11.4 to 10.2 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 03
WINDII = 01
HALOE = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 01
Two (2) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
One (1) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT finished testing the Eclipse software patch to fix erroneous
alarm messages. It is now installed on the realtime system and has
been working successfully since installation. The Raytheon
development team has also solved the problem seen on the status buffer
display and will deliver a new database this week for testing.
MEETINGS
-None this report period
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 02 August 2002 (DOY 214).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT
Occurrences: Appendix C, Clock & Solar Array Rate
Corrections: Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 26 July 2002
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
59375 201/2200 02-254 WINDII POWER OFF
59377 202/0056 02-255 OBC/CLK CLK ADJ FR 0.00 TO 20.6 MSEC/DAY
59385 202/1335 02-256 HALOE POWER ON
59397 203/0817 02-257 OBC ATC NOOP;XMTR-ON CMD
59400 203/1334 02-258 OBC/CLK CLK ADJ FR 20.6 TO 0.00 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 26 July 2002
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
______________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 26 July 2002
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
59377 202/0103 CLK ADJ FR 0.00 TO 20.6 MSEC/DAY
59400 203/1339 CLK ADJ FR 20.6 TO 0.00 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 26 July 2002
19 July 2002 - GMT Day 200 Beta = 58.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -157.0/-123.7 -3.30/-4.52 +00.0/+00.0 8.4 0.000 0.0
2 + 11.2/- 67.2 +3.87/+2.60 +20.8/-13.5 25.8 1.030 17.7
3 + 22.4/- 33.6 +2.92/+1.97 +20.3/-13.4 25.6 1.020 17.3
26 July 2002 - GMT Day 207 Beta = 28.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -157.0/-168.0 -3.30/-4.24 +00.0/+00.0 08.5 0.000 0.0
2 + 11.2/- 69.6 +6.44/+5.15 +27.5/-17.9 27.8 1.030 21.8
3 + 22.4/- 56.0 +5.21/+3.87 +26.7/-17.6 27.7 1.020 21.4
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period
APPENDIX E
UARS ANOMALY REPORTS
Week ending 26 July 2002
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-039 FOT 59368 201/1002 30832 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-039 FOT See List July 30832 Generic Late Acquisition
02-041 FOT 59279 195/1156 N/A TMON 01 tripped